# Incorporating Gradients to Rules: Towards Lightweight, Adaptive Provenance-based Intrusion Detection

### v1

@@ Abstract
As cyber-attacks become increasingly sophisticated and stealthy, it becomes more imperative and challenging to detect intrusion from normal behaviors. Through fine-grained causality analysis, provenance-based intrusion detection systems (PIDS) demonstrated a promising capacity to distinguish benign and malicious behaviors, attracting widespread attention from both industry and academia. Among diverse approaches, rule-based PIDS stands out due to its lightweight overhead, real-time capabilities, and explainability. However, existing rule-based systems suffer low detection accuracy, especially the high false alarms, due to the lack of fine-grained rules and environment-specific configurations.
In this paper, we propose Captain, a rule-based PIDS capable of automatically adapting to diverse environments. Specifically, we propose three adaptive parameters to adjust the detection configuration with respect to nodes, edges, and alarm generation thresholds. We build a differentiable tag propagation framework and utilize the gradient descent algorithm to optimize these adaptive parameters based on the training data. We evaluate our system based on data from DARPA Engagement and simulated environments. The evaluation results demonstrate that Captain offers better detection accuracy, less detection latency, lower runtime overhead, and more interpretable detection alarms and knowledge compared to the SOTA PIDS.
@@ I Introduction
Advanced Persistent Threats (APT) are becoming a growing threat to both government and industrial sectors, causing significant societal impacts [3]. The Equifax breach in 2017 resulted in the theft of vast amounts of personal data, highlighting the severe privacy and security risks [35, 42]. Moreover, attackers continuously innovate to find new ways to penetrate systems and remain undetected for extended periods. Recently, provenance-based intrusion detection systems (PIDS) have gained attention from both the security industry and academia for their causality analysis capability. However, prevailing alarm fatigue, excessive runtime overhead, and long detection latency are still open research problems in real-world scenarios [13, 38, 7, 17].
Alarm fatigue [12, 39] is a significant issue plaguing the security industry. A recent survey [1] indicates that security analysts are required to handle an overwhelming average of 5,000 alarms daily, with a majority being false alarms. The excessive volume of alarms can lead to serious consequences. For instance, in the 2013 Target data breach [35], the malicious activities were detected and reported by security tools but overlooked by analysts, resulting in delayed response and expanded losses [11]. In practice, false alarms are equally detrimental as missed alarms. Furthermore, the time needed for threat investigation largely depends on the detection granularity and attack semantics the alarms provide, such as which system entities and events form the attack chain and how they are interconnected. The overload and ambiguity of alarms make it hard for analysts to handle these alarms effectively.
In practice, significant challenges also arise from runtime overhead and detection latency. High runtime overhead compromises the scalability of detection systems, hindering their deployment on a large scale [13]. Moreover, detection latency critically affects performance, as prolonged latency delays the response to threats, thereby impeding analysts’ efficiency in managing alarms.
In academia, a recent trend of PIDS [51, 48, 9, 34] leverage embedding techniques like word2vec [34] and graph2vec [48] to encode system entities and events, and neural networks like Graph Neural Networks (GNNs) [34, 9, 51] and Recurrent Convolutional Neural Networks (RCNNs) [48] to analyze information flow and produce detection results. We refer to them as embedding-based PIDS. Although these systems have demonstrated notable achievements in detection performance, they face the following challenges:
@@
On the other hand, a group of PIDS [18, 19, 30, 46, 17] operates based on human-defined mechanisms to generate representations for system entities, propagate information flow, and trigger alarms. We refer to them as rule-based PIDS. Leveraging these rules, they take actionable leads, allowing themselves to provide semantic-rich alarms and pinpoint specific suspicious events for further investigation. Besides, the rule-based approach operates on simple arithmetic calculations, which require far fewer computational resources than complex matrix calculations by neural networks. Furthermore, some rule-based approaches can consume streaming data once it arrives, which greatly reduces storage requirements and allows fast response with minimal delay. Rule-based PIDS shows advantages in real-world detection tasks due to its fine-grained detection, semantic-rich alarms, lightweight overhead, and minimal detection delay.
@@
Despite rule-based PIDS excelling in the aforementioned perspectives, it also faces challenges in real-world deployment. Existing rule-based PIDS suffer from the inflexibility to adapt to different environments, thus having a high volume of false alarms during detection. In SOCs, analysts have to manually configure the models, which is a time-consuming process [22]. This highlights the need for an automated configuration methodology and scheme to enhance the accuracy of the rule-based PIDS. To achieve this, we require a rule-based PIDS to adjust its configuration based on the detection feedback on the training set without human interference. Thus, an automatic feedback mechanism is needed. This necessitates establishing a linkage between the configuration and the detection results, enabling the configuration to be dynamically adjusted based on the outcomes of the detection process.
@@
In this paper, we propose Captain, a rule-based PIDS with adaptive configuration learning capability. Our work aims to leverage the advantages of traditional rule-based PIDS while enabling the system to acquire suitable configurations autonomously. Specifically, we introduce three adaptive parameters to the rule-based PIDS, giving it more flexibility during detection. We construct a learning module for these parameters. We calculate and record the gradients of each adaptive parameter, transferring the rule-based system to a differentiable function. With the help of a loss function, the gradient descent algorithm can be utilized to optimize the adaptive parameters based on the training data and thus reduce the false alarms in the testing phase.
@@
We evaluate Captain in diverse benign and APT attack scenarios, which are drawn from the widely acknowledged public datasets [5] provided by DARPA Engagement [33] and datasets collected from simulated environments in collaboration with a SOC. The evaluation results demonstrate that Captain can reduce false alarms by over 90% (11.49x) on average compared to traditional rule-based PIDS. Additionally, Captain achieves much better detection accuracy with less than 10% CPU usage and significantly lower memory usage and latency compared to the SOTA embedding-based PIDS. We then select a few scenarios as cases to study the explainability of the learned configurations.
All in all, this paper makes the following contributions:
@@
  • We designed Captain, a PIDS that is based on general tag-propagation rules and can adjust its rules according to normal data to reduce false alarms. This design endows Captain with the lightweight, low latency, interpretability, and fine detection granularity of rule-based PIDS while also having excellent adaptability as a learning-based PIDS.
@@
  • We parameterize the rule-based system and propose a novel differentiable tag propagation framework. In doing so, the gradient descent algorithm can be utilized to obtain the optimal parameters in different detection environments.
@@
  • We systemically evaluate Captain in various scenarios, including commonly used DARPA datasets and simulated datasets from a partner SOC. The evaluation result demonstrates that Captain can automatically adapt to various environments and provide better detection accuracy with less latency, lower overhead, and more interpretable results.
@@ II Background and Related Work
In this section, we will first introduce PIDS and highlight mainstream approaches to elucidate our design choice. Then, we introduce some background about the optimization problem, as well as describe our threat model and assumptions.
@@ II-A Provenance-based Intrusion Detection
Recently, provenance-based intrusion detection has drawn wide attention from industry and academia for its powerful correlation and causal analysis capabilities. When performing provenance analysis, the system behaviors are modeled as directed acyclic graphs, called provenance graphs [52, 26], in which nodes represent system entities, such as processes, files, sockets, pipes, memory objects, etc, and edges represent interactions between entities, such as reading files and connecting to a remote host.
@@
Fig. 1 illustrates an overview of typical PIDS workflow. PIDS exploits the graph nature of provenance data for whole attack information flow (data flow and control flow) analysis or multi-stage attack technique correlation analysis.
@@
Based on the way to model and aggregate the entity information, the data/control flow, and the system behaviors, we identify two major groups of PIDS: the embedding-based approach and the rule-based approach.
@@
Embedding-based PIDS. Embedding-based systems [25, 16, 50, 2, 51, 34, 9] use numerical vectors to represent, propagate, and aggregate the information of system entities, events, data/control flow in the graph. Various techniques in machine learning have been employed to embed the graph features into numerical vectors. Since these techniques typically require the input structured as graphs or paths, embedding-based PIDS usually adopt time windows or log batches in their design leading to the additional detection delay, which we named as buffer time in this paper.
Concretely, buffer time refers to the duration of the time window in which all streaming logs are stored to create a graph structure for subsequent processing. Previous studies have implemented a fixed-length window based on time [9] or the number of events [34], while some works [48] utilize the variable length window. Buffer time is crucial in real-world detection scenarios because, in the worst-case scenario, after an attack has been launched, the PIDS would have to wait for the duration of the buffer time before starting the detection procedure.
@@
For instance, ProvDetector [43] selects rare paths based on historical event frequency to generate embedding vectors for sequence learning and anomaly detection, which requires the entire path to perform detection. Unicorn[16] embeds the extracted histograms from the provenance graph and clusters the embedded graph sketches. Atlas [2] extracts sequences from the graph and learns to classify them with Long Short-term Memory (LSTM). Watson [50] learns the node embedding and relative event importance from the benign traces. More recent embedding-based systems employed graph neural networks [34, 9] to detect threats. While such systems can achieve better detection accuracy, they fail to provide the rationale for generated alarms. For example, ShadeWatcher [51] provides the probability of malicious system events without further context. Flash [34] highlights suspicious entities with causal links but does not provide further attack semantics of those interactions. In practice, more manual efforts are required to review the relevant audit logs thoroughly for further response.
@@
Rule-based PIDS. Rule-based systems [18, 46, 30, 19, 17] leverage human-designed rules to map system entities and events to predefined semantic units (e.g., tags [18, 19, 46], TTP [30], etc.) and propagation these units along the information flow to carry out the detection results. NoDoze[17] assigns an anomaly score to each edge in the provenance graph based on the frequency it appears in the benign data and propagates the anomaly scores on the graph. Poirot [29] models APT detection as a graph pattern matching problem that aligns manually constructed query graphs on the provenance graph. Rule-based systems have become popular due to their real-time processing capabilities and lightweight properties. However, they usually suffer low detection accuracy due to the lack of fine-grained rules and environment-specific configurations.
@@
Table I summarizes the detection granularity and whether performing detection requires buffer time.
@@
In this paper, we build our system using the rule-based detection method for the following reasons: 1) rule-based detection is computationally more efficient [29, 16, 18]; 2) rule-based methods are suitable for real-time detection because they can process the event stream incrementally [26]; 3) rule-based methods [18, 19, 46] offer more explainability to the detection process and results than embedding-based methods [51, 17, 44, 16, 34, 9]; 4) rule-based methods are robust against mimicry attacks [15].
@@
1Although Kairos triggers alarms on the graph level, it can highlight anomaly nodes and edges in the subgraphs
@@ II-B Configuration as an Optimization Problem
Finding the best configuration for a rule-based PIDS can be formalized as an optimization problem, which, in short, aims to find the best elements $\theta$ in a searching space to minimize or maximize an objective function $J(\theta)$. Gradient descent is one of the most common algorithms to solve the optimization problem [37]. It utilizes the gradient of the objective function to the parameters $\nabla_{\theta}J(\theta)$ to update $\theta$ according to $\theta=\theta-l\cdot\nabla_{\theta}J(\theta)$, where $l$ is the learning rate. Many variants have been proposed based on the gradient descent algorithm, including stochastic gradient descent [36], mini-batch gradient descent, Adagrad [14], Adam [21], and so on. We discuss more algorithms and their applicability to Captain in § VII-A.
While the optimization problem and its gradient-based solutions play a significant role in quantitative science and engineering [28], it is rarely discussed in the previous rule-based PIDS because of the following challenges. First, no existing work formally defined the adjustable parameters with clear meanings in the context of rule-based PIDS. Second, designing an objective function, calculating, updating, and saving the gradients needed in the gradient descent algorithm are knotty problems for a tag-based PIDS. In this paper, we introduce Captain to address these two identified gaps.
@@ II-C Threat Model & Assumptions
Similar to many previous works [19, 18, 16, 44, 17, 25], we assume that OS kernels and auditing tools are a part of the trusted computing base (TCB), which means the attackers are not able to tamper system auditing data. Meanwhile, we do not take hardware manipulation and any other attacks that leave no traces in terms of the provenance graph into consideration.
@@ III Motivation Examples
In this section, we use three motivating examples to elucidate three common challenges encountered by rule-based PIDS, as depicted in Fig. 2.
Figure 2: Motivating examples from the Engagement dataset. (a) Communication between different IP addresses and sshd process. (b) Download activities via wget. (c) Two kinds of bash script execution we observed.
Entity-level Distinction: Entity-level information plays a pivotal role in understanding system behaviors through provenance graphs. System entities span various categories, such as processes, files, sockets, etc. Each system entity is characterized by specific features: processes are identified by their names and command lines, files by their file paths, and sockets by their IP addresses and port numbers.
Embedding-based PIDS usually use machine learning techniques, especially from the natural language processing (NLP) field, to embed the entity-level features, such as hierarchical feature hashing [9], Word2Vec [34], and FastText [25]. However, these techniques are computationally expensive and inexplicable. On the other hand, it is hard for rule-based PIDS to distinguish the system entities for the following reasons: 1. the amount of entities is huge; 2. the detailed entity-level configuration rules vary across different environments. Therefore, in most cases, they have to apply general rules. Unfortunately, such generic configurations often result in excessive false alarms. As the motivating example in Fig. 2(a) shows, if the detection systems do not distinguish the network nodes by their IP addresses, all system entities interacting with sshd would inherit maliciousness from those external connections and trigger false alarms.
Event-level Distinction: Event-level distinction is also challenging for the rule-based PIDS during detection. System events serve as bridges that propagate the contextual information flows among system entities and form the kill chains for the attack. However, traditional rule-based PIDS find it challenging to adaptively generate fine-grained propagation and detection rules for different environments. Generic rules might lead to unsatisfying detection performance. In our motivating example in Fig. 2(c), firefox and sshd are two processes that often have network activities with various IP addresses. However, sshd is more likely to use bash to execute scripts, while it is suspicious for firefox to perform such tasks. Please note that in this example, the related entities and event types are the same for firefox and sshd. To distinguish such two behaviors, a very explicit rule is needed (A knotty question would be: do we need to set different rules about the “execute bash” event for each different process?). Some rule-based PIDS used frequency information in the training set to generate rules [17]. We will show later that their methods are more susceptible to the training set poisoning attack.
Dependency Explosion: Besides distinguishing system entities and events, dependency explosion is a common problem in provenance analysis [19]. In our motivating example in Fig. 2(b), we observe that the bash processes frequently write to and read from the .bash_history file. Notably, the bash process interacts with several potentially malicious entities as demonstrated in Fig. 2(a) and (c). Therefore, the bash processes tend to have relatively high maliciousness in these scenarios. Consequently, these bash processes propagate the maliciousness to the .bash_history file and then .bash_history spread the maliciousness among other system entities via subsequent interactions, triggering more false alarms. Many existing efforts have been made to solve this issue [32, 47, 4, 27, 24]. Unfortunately, their solutions usually require extensive instrumentation of applications/OS [4, 27, 24] and expertise, which limits the usage in the real-world system, especially in end-point host deployment. In this paper, we aim to address the dependency explosion issue via automatic learning without instrumentation on the auditing module.
IV Captain Design

### v2

@@ Abstract
As cyber attacks grow increasingly sophisticated and stealthy, it becomes more imperative and challenging to detect intrusion from normal behaviors. Through fine-grained causality analysis, provenance-based intrusion detection systems (PIDS) demonstrated a promising capacity to distinguish benign and malicious behaviors, attracting widespread attention from both industry and academia. Among diverse approaches, rule-based PIDS stands out due to its lightweight overhead, real-time capabilities, and explainability. However, existing rule-based systems suffer low detection accuracy, especially the high false alarms, due to the lack of fine-grained rules and environment-specific configurations.
In this paper, we propose Captain, a rule-based PIDS capable of automatically adapting to diverse environments. Specifically, we propose three adaptive parameters to adjust the detection configuration with respect to nodes, edges, and alarm generation thresholds. We build a differentiable tag propagation framework and utilize the gradient descent algorithm to optimize these adaptive parameters based on the training data. We evaluate our system using data from DARPA Engagements and simulated environments. The evaluation results demonstrate that Captain enhances rule-based PIDS with learning capabilities, resulting in improved detection accuracy, reduced detection latency, lower runtime overhead, and more interpretable detection procedures and results compared to the state-of-the-art (SOTA) PIDS.
Note: This is a preprint version of the paper accepted at the Network and Distributed System Security Symposium (NDSS) 2025.
@@ I Introduction
Advanced Persistent Threats (APT) are becoming a growing threat to both government and industrial sectors, causing significant societal impacts [1]. The Equifax breach in 2017 resulted in the theft of vast amounts of personal data, highlighting the severe privacy and security risks [2, 3]. Moreover, attackers continuously innovate to find new ways to penetrate systems and remain undetected for extended periods. In recent years, provenance-based intrusion detection systems (PIDS) have gained attention from both the security industry and academia for their causality analysis capability. However, prevailing alarm fatigue, excessive runtime overhead, long detection latency, and opaque detection processes (based on black-box) are still open research problems in real-world scenarios [4, 5, 6, 7, 8].
Alarm fatigue [9, 10] is a significant issue plaguing the security industry. A recent survey [11] indicates that security analysts are required to handle an overwhelming average of 5,000 alarms daily, with a majority being false alarms. The excessive volume of alarms can lead to serious consequences. For instance, in the 2013 Target data breach [2], the malicious activities were detected and reported by security tools but overlooked by analysts, resulting in delayed response and expanded losses [12]. In practice, false alarms are equally detrimental as missed alarms. Moreover, significant challenges also arise from runtime overhead and detection latency. High runtime overhead compromises the scalability of detection systems, hindering their deployment on a large scale [4]. Moreover, detection latency critically affects performance, as prolonged latency delays the response to threats, thereby impeding analysts’ efficiency in managing alarms.
In academia, a recent trend of PIDS [13, 14, 15, 16] leverage embedding techniques like word2vec [16] and graph2vec [14] to encode system entities and events, and neural networks like Graph Neural Networks (GNNs) [16, 15, 13, 17] and Recurrent Convolutional Neural Networks (RCNNs) [14] to analyze information flow and produce detection results. We refer to them as embedding-based PIDS. Although these systems have demonstrated notable achievements in detection performance, they face the following challenges:
@@
On the other hand, a group of PIDS [21, 22, 23, 24, 7] operates based on human-defined mechanisms to generate representations for system entities, propagate information flow, and trigger alarms. We refer to them as rule-based PIDS. Leveraging these rules, they take actionable leads, allowing themselves to provide semantic-rich alarms and pinpoint specific suspicious events for further investigation. Besides, the rule-based approach operates on simple arithmetic calculations, which require far fewer computational resources than complex matrix calculations by neural networks. Furthermore, some rule-based approaches process streaming data once it arrives, significantly reducing storage requirements and enabling rapid response with minimal delay. In summary, rule-based PIDS shows advantages in real-world detection tasks due to its fine-grained detection, semantic-rich alarms, lightweight overhead, and minimal detection delay.
@@
Despite rule-based PIDS excelling in the aforementioned perspectives, they also face challenges in real-world deployment. These systems often employ simplistic and universal rules to distinguish between graphs (such as discrete trustworthiness levels of node in [22, 21, 23], universal system parameters in [22, 21, 23, 7]). The inflexible rules lead these PIDS to be either too lenient or too strict. In security operations centers (SOCs), analysts have to manually configure the models, which is a time-consuming process [25]. This highlights the need for an automated configuration methodology and scheme. To achieve this, a rule-based PIDS must be capable of adjusting its rules autonomously based on detection feedback from the training set, without human intervention. This requires an automatic feedback mechanism that establishes a direct connection between the rules and detection results, allowing the system to dynamically adjust its configuration based on the detection outcomes.
@@
In this paper, we propose Captain, a rule-based PIDS with adaptive configuration learning capability. We aim to leverage the advantages of traditional rule-based PIDS while enabling the system to acquire suitable configurations autonomously. Specifically, we introduce three adaptive parameters to the rule-based PIDS, giving it more flexibility during detection. Moreover, we design a learning module to adjust these parameters automatically based on the detection results during training. We calculate and record the gradients of each adaptive parameter, transferring the rule-based system to a differentiable function. With the help of a loss function, the gradient descent algorithm can be utilized to optimize the adaptive parameters based on the benign training data and thus reduce the false alarms in testing.
@@
We evaluate Captain in diverse detection scenarios, which are drawn from the widely acknowledged public datasets [26] provided by DARPA Engagement [27] and datasets collected from simulated environments in collaboration with a SOC. The evaluation results demonstrate that Captain can reduce false alarms by over 90% (11.49x) on average compared to traditional rule-based PIDS. Additionally, Captain achieves much better detection accuracy with less than 10% CPU usage and significantly lower memory usage and latency compared to the SOTA embedding-based PIDS. We then select a few scenarios as cases to study the explainability of the learned configurations.
In summary, this paper makes the following contributions:
@@
  • We propose Captain, a rule-based PIDS that can adjust its rules using benign data to reduce false alarms. Our design endows Captain with the lightweight nature, low latency, and interpretability of rule-based PIDS while also having fine detection capability and adaptability.
@@
  • We parameterize the rule-based PIDS and propose a novel differentiable tag propagation framework that allows us to optimize adaptive parameters using the gradient descent algorithm.
@@
  • We systemically evaluate Captain across various scenarios, including widely used DARPA datasets and simulated datasets from a partner SOC. The evaluation result demonstrates that Captain can automatically adapt to various environments, offering superior detection accuracy with reduced latency, lower overhead, and more interpretable detection.
@@ II Background and Related Work
In this section, we begin by introducing PIDS and highlighting mainstream approaches to clarify our design choices. Next, we provide background on the optimization problem, followed by a description of the threat model and the assumptions.
@@ II-A Provenance-based Intrusion Detection
Provenance-based intrusion detection has drawn wide attention for its powerful correlation and causal analysis capabilities. In provenance analysis, the system behaviors are modeled as directed acyclic graphs, called provenance graphs [28, 29], in which nodes represent system entities, such as processes, files, sockets, pipes, memory objects, etc., and edges represent interactions between entities, such as reading files and connecting to a remote host.
@@
Fig. 1 illustrates an overview of typical PIDS workflow. Based on how to model and aggregate the entity information, the data/control flow, and the system behaviors, we identify two major groups of PIDS: the embedding-based approach and the rule-based approach.
@@
Embedding-based PIDS. Embedding-based systems [18, 19, 30, 31, 13, 16, 15, 17] use numerical vectors to represent, propagate, and aggregate the information of system entities, events, data/control flow in the graph. Various techniques in machine learning have been employed to embed the graph features into numerical vectors. Since these techniques typically require the input structured as graphs or paths, embedding-based PIDS usually adopt time windows or log batches in their design, leading to the additional detection delay, which we named as buffer time in this paper. Concretely, buffer time refers to the time window duration in which all streaming logs are stored to create a graph structure for subsequent processing. Previous studies have implemented a fixed-length window based on time [15] or the number of events [16], while some works [14] utilize the variable length window. Buffer time is crucial in real-world detection scenarios because, in the worst-case scenario, after an attack has been launched, the PIDS would have to wait for the duration of the buffer time before starting the detection procedure.
@@
The methods for embedding graphs differ across various systems. For instance, ProvDetector [32] selects rare paths based on historical event frequency to generate embedding vectors for sequence learning and anomaly detection. Atlas [31] classifies the extracted sequences from the graph with Long Short-term Memory (LSTM). Unicorn[19] embeds the graph histograms and clusters the embedded graph sketches. Watson [30] learns the node embedding and event semantics from training traces. More recent embedding-based systems [16, 15, 17] employed graph neural networks in detection. While such systems can achieve better detection accuracy, they fail to provide the rationale for generated alarms. For example, ShadeWatcher [13] provides the probability of malicious system events without further context. Flash [16] highlights suspicious entities with causal links but does not provide further attack semantics of those interactions. In practice, more manual efforts are required to review the relevant audit logs thoroughly for further response.
@@
Rule-based PIDS. Rule-based systems [21, 24, 23, 22, 7] leverage human-designed rules to map system entities and events to predefined semantic units (e.g., tags [21, 22, 24], TTP [23], etc.) and propagation these units along the information flow to capture causality. NoDoze[7] assigns an anomaly score to each edge based on the frequency it appears in the benign data and propagates the scores on the graph. Poirot [33] models the detection as a graph-matching problem that aligns manually constructed query graphs on the provenance graph. Sleuth [21] and Morse [22] assign tags to system entities and propagate tags among the graph. Rule-based systems have become popular due to their real-time processing capabilities and lightweight properties. However, they usually suffer low detection accuracy due to the lack of fine-grained rules and environment-specific configurations. Most rule-based PIDS [22, 21, 24, 23] implement simple and universal rules, making them inflexible to adapt to different environments. These systems typically rely on a discrete classification to denote system entity properties, such as private/public, and trusted/untrusted, and apply identical rules to all nodes and events. As a result, they lack the nuance required to accurately distinguish between similar graphs, often being either overly lenient or overly strict, which leads to false alarms or missed true alarms. We will elaborate on these challenges in §III.
@@
Table I summarizes the detection granularity of existing PIDS and indicates whether the detection requires buffer time.
@@
In this paper, we build our system Captain following the rule-based methodology for four reasons: 1) rule-based detection is computationally more efficient [33, 19, 21]; 2) rule-based methods are suitable for real-time detection because they can process the event stream incrementally [29]; 3) rule-based methods offer more explainability to the detection process and results than embedding-based methods [34]; 4) rule-based methods are robust against mimicry attacks [35]. Additionally, Captain introduces a significant advancement over traditional rule-based PIDS. Unlike systems that rely on simple, universal, and inflexible rules, Captain offers a mathematically complete framework for automatically fine-tuning detection rules, enabling more precise detection and response to security threats.
@@
1Although Kairos triggers alarms on the graph level, it can highlight anomaly nodes and edges in the subgraphs.
@@ II-B Configuration as an Optimization Problem
Finding the best configuration for a rule-based PIDS can be formalized as an optimization problem, which, in short, aims to find the best elements $\theta$ in a searching space to minimize or maximize an objective function $J(\theta)$. Gradient descent is one of the most common algorithms to solve the optimization problem [36]. It utilizes the gradient of the objective function to the parameters $\nabla_{\theta}J(\theta)$ to update $\theta$ according to $\theta=\theta-l\cdot\nabla_{\theta}J(\theta)$, where $l$ is the learning rate. Many variants have been proposed based on the gradient descent algorithm. We discuss more algorithms and their applicability to Captain in § VII-B.
While the optimization problem and its gradient-based solutions play a significant role in quantitative science and engineering [37], it is rarely discussed in the previous rule-based PIDS because of the following challenges. First, no existing work formally defined the adjustable parameters with clear meanings in the context of rule-based PIDS. Second, designing an objective function, calculating, updating, and saving the gradients needed in the gradient descent algorithm are knotty problems for a rule-based PIDS.
@@ II-C Threat Model & Assumptions
Similar to many previous works [22, 21, 19, 32, 7, 18], we assume that OS kernels and auditing tools are a part of the trusted computing base (TCB), which means the attackers are not able to tamper system auditing data. Meanwhile, we do not consider hardware manipulation and any other attacks that leave no traces on the provenance graph.
@@ III Motivation Examples
In this section, we present motivating examples to elucidate the challenges encountered by traditional, coarse-grained rule-based PIDS, as depicted in Fig. 2. These scenarios contain similar benign and malicious behaviors, which pose challenges to existing rule-based PIDS [22, 21, 23, 7] in adjusting their rules to effectively select an optimal trade-off point between excessive false alarms and missing true alarms.
Figure 2: Three motivating examples from the real-world dataset where the more fine-grained rules are needed in the rule-based PIDS.
“Grey” Nodes: With the advancement of cloud computing, connections to Function as a Service (FaaS) platforms like AWS Lambda and Cloudflare Workers have become a typical pattern in provenance graphs. However, recent reports indicate that attackers are exploiting these commercial FaaS services to redirect traffic to Command and Control (C2) servers, as illustrated in Fig. 2a. Existing rule-based PIDS are often too coarse to depict those IPs. For example, Sleuth [21] only defined three trustworthiness levels: Benign Authentic, Benign, and Unknown for the nodes, and Morse [22] used a binary initial tag for the trustworthiness. But in this scenario, since both benign and malicious behavior can involve communications with the external FaaS IPs, the coarse-grained rules either lead to excessive false alarms or missing true alarms. A more fine-grained configuration to reflect the “grey scale” of these IPs may help solve the issue, i.e., we can assign a trustworthiness value (a float between 0 and 1) to every node in the provenance graph.
Dependency Explosion: Dependency explosion refers to the scenario where all the subsequent events are treated as they depend on the previous ones, and thus a single suspicious event can lead to millions of system entities considered suspicious [22]. As shown in Fig. 2b, a prolonged and furcate event chain can spread the maliciousness carried by the suspicious entities to entities involved in all subsequent events. Many existing efforts have been made to solve this issue [38, 39, 40, 41, 42]. Unfortunately, their solutions usually require extensive instrumentation of applications/OS [40, 41, 42] and expertise, which limits the usage in the real-world system, especially in end-point host deployment.
To deal with the dependency explosion problem without extensive instrumentation, existing rule-based PIDS adopt methods like event whitelisting [21], cost-based pruning, or dividing suspicious events into different stages in a typical APT lifecycle [23]. Morse [22] proposed the decay and attenuation mechanisms to reduce the maliciousness carried by the system entities over time to mitigate the dependency explosion. However, the decay and attenuation factor is the same for all nodes and edges. Such identical configuration can lead to the dilemma between excessive false alarms and missing true alarms as attackers may deliberately fork many irrelevant processes or connect to benign entities to evade detection. As shown by the two provenance graphs in Fig. 2b, we want to confine the dependency explosion in the left graph while preserving the maliciousness in the right (the right graph depicts a rare pattern where a firefox process forks a Python process, indicating the possibility of suspicious script execution from a website.) Ideally, we aim to set a propagation rate for each individual event, allowing for more customized control. This ensures that the maliciousness of unusual events does not fade away too quickly, while it can reset rapidly for more common events.
Customized Alarm Triggering: In rule-based PIDS, alarms are triggered based on predefined conditions that are uniformly applied to all events, regardless of their context. These one-fits-all approaches can lead to suboptimal detection performance. For instance, Morse sets a series of detection rules and thresholds to trigger alarms and the rules are identical for all processes and events. As the motivating example shows in Fig. 2c, firefox and sshd are two processes that often have network activities with various IP addresses. While sshd may commonly use bash for script execution, it is unusual and suspicious for firefox to do the same. Despite the similarities in event types and related entities between firefox and sshd, applying a uniform threshold for all “fork” events could either result in false alarms for sshd or miss genuine threats involving firefox. To distinguish between such two behaviors, a specific rule is needed: for the common events, we want to set a higher threshold to avoid excessive false alarms, while for rare events, we want to keep the threshold low to avoid missing true alarms. A knottier question then follows: how to set different thresholds about the “execute bash” event for each process?
The three motivating examples underscore the challenges that existing rule-based PIDS face in balancing false and true alarms due to their coarse-grained rules. Additionally, these systems lack the flexibility to dynamically adjust rules to accommodate different environments. Theoretically, embedding-based PIDS, leveraging the power of deep learning, can effectively create customized “rules” for every event and entity in the graph by embedding them into the feature space. However, training and deploying deep neural networks require numerous calculation resources and storage space, leading to high runtime overhead. Moreover, these PIDS usually lack the explainability of the detection process and result, making it difficult to pinpoint the malicious behaviors at the event level.
To overcome these limitations, we aim to refine rule-based PIDS using fine-grained rules. By allowing the rule-based detector to learn and adapt from training data, we can enhance detection accuracy without the extensive resource demands and runtime overhead. Our system improves the overall efficacy of rule-based PIDS but also maintains its operational simplicity and clarity.
IV System Design